### v2

# Weak metric structures on generalized Riemannian manifolds

Abstract.
Linear connections with torsion are important in the study of generalized Riemannian manifolds $(M,G=g+F)$, where the symmetric part $g$ of $G$ is a non-degenerate (0,2)-tensor and $F$ is the skew-symmetric part. Some space-time models in theoretical physics are based on $(M,G=g+F)$, where $F$ is defined using an almost complex or almost contact metric structure.
In the paper, we first study more general models, where $F$ has constant rank and is based on weak metric structures (introduced by the first author and R. Wolak), which generalize almost complex and almost contact metric structures. We consider generalized metric connections (i.e., linear connections preserving $G$) with totally skew-symmetric torsion (0,3)-tensor. For rank$(F)=\dim M$ and non-conformal tensor $A^{2}$, where $A$ is a skew-symmetric (1,1)-tensor adjoint to $F$, we apply weak almost Hermitian structures to fundamental results (by the second author and S. Ivanov) on generalized Riemannian manifolds and prove that the manifold is a weighted product of several nearly Kähler manifolds corresponding to eigen-distributions of $A^{2}$. For rank$(F)<\dim M$ we apply weak $f$-structures and obtain splitting results for generalized Riemannian manifolds.
Keywords: weak almost Hermitian manifold, weak metric $f$-structure, metric connection, totally skew-symmetric torsion, generalized Riemannian manifold
Mathematics Subject Classifications (2010) 53C15, 53C25, 53B05.
1. Introduction
The metric-affine geometry (founded by E. Cartan) generalizes Riemannian geometry: it uses a linear connection with torsion, $\nabla$, instead of the Levi-Civita connection of a Riemannian metric $g$, and appears in such context as homogeneous and almost Hermitian manifolds, Finsler and generalized geometries, e.g. [21]. The important distinguished case is Riemann-Cartan manifolds, where metric connections ($\nabla g=0$) are used. Riemann-Cartan spaces are central in gauge theory of gravity, where the torsion tensor is involved in the Cartan spin-connection equation.
Starting from 1950, A. Einstein used the real non-symmetric basic tensor $G$, sometimes called generalized Riemannian metric/manifold. In this theory the symmetric part $g$ of the basic tensor $G=g+F$ is related to gravitation, and the skew-symmetric one $F$ to electromagnetism.
In this paper, we investigate linear connections on a generalized Riemannian manifold $(M,G)$ with a general assumption, that the symmetric part $g$ of $G$ is non-degenerate. We extend examples of generalized Riemannian manifolds by weak almost (para) Hermitian manifolds, weak (para) $f$-manifolds and weak almost (para) contact manifolds (defined by V. Rovenski and R. Wolak [24]). We consider linear connections with torsion preserving the non-symmetric tensor $G$, so it means that it preserves its symmetric and skew-symmetric parts, $\nabla g=\nabla F=0$, and consequently it preserves $Q$, $\nabla Q=0$, where $Q$ is a self-adjoint endomorphism with respect to $g$ and $Q$ takes different values depending on the various types of manifolds mentioned above.
In many geometric and physical contexts, such as string theory, gravity, and complex or almost complex geometry, skew-symmetric tensor $F$ plays a fundamental role [3, 10, 12, 18, 27]. Generalized Riemanian manifolds studied here allow the natural incorporation of metric connections with torsion, particularly those preserving both $g$ and $F$, and provide a unified framework for exploring richer geometric structures beyond the scope of (pseudo-) Riemannian geometry.
We pay special attention to the case when the torsion of the connection preserving the generalized Riemannian metric $G$ is totally skew-symmetric with respect to the symmetric part $g$ of $G$ and satisfies the $A$-torsion or $Q$-torsion condition, depending on the structure. The corresponding new Nijenhuis tensors naturally appear in weak almost (para-)Hermitian, and weak almost (para-)contact manifolds, whose total skew-symmetry depends on the $A$-torsion condition or on the $Q$-torsion condition. A Codazzi connection is defined on generalized Riemannian manifolds, and an example of such a connection is given. The Chern and Bismut connections are also studied in the context of weak Hermitian structures.
The paper is organized as follows. In Section 2, we consider the geometric model. We introduce Codazzi connections and metric connections adapted to generalized Riemannian manifolds, and discuss the torsion and Nijenhuis tensors, including their properties and significance. Section 3 presents applications of weak metric structures: weak almost Hermitian structures, weak $f$-structures satisfying $f^{3}+fQ=0$, weak almost contact metric structures, and weak para-Hermitian and weak para-$f$-structures with $f^{3}-fQ=0$, to generalized Riemannian manifolds.
2. Linear Connections on Generalized Riemannian Manifolds
2.1. Geometrical Model
The fundamental (0,2)-tensor $G$ in a non-symmetric (generalized) Riemannian manifold $(M,G)$ is in general non-symmetric. It is decomposed in two parts, the symmetric part $g$ (called Riemannian metric) and the skew-symmetric part $F$ (called fundamental 2-form), $G=g+F$, where
$g(X,Y)=\frac{1}{2}\big{[}G(X,Y)+G(Y,X)\big{]},\qquad F(X,Y)=\frac{1}{2}\big{[}G(X,Y)-G(Y,X)\big{]}.$
We assume that the symmetric part, $g$, is non-degenerate of arbitrary signature, and the skew-symmetric part, $F$, has a constant rank, e.g., is non-degenerate. Therefore, we obtain a well-defined (1,1)-tensor $A$ of constant rank determined by the following condition:
(2.1) $g(AX,Y)=F(X,Y)\quad\text{for all }X,Y\in\mathfrak{X}_{M}.$
By the above, since $F$ is skew-symmetric, then the tensor $A$ is also skew-symmetric: $g(AX,Y)=-g(X,AY)$ for all $X,Y\in\mathfrak{X}_{M}$. We seek natural linear connections $\nabla$ on $M$ preserving the generalized Riemannian metric $G,\ \nabla G=0$, with a torsion (1,2)-tensor $T(X,Y)=\nabla_{X}Y-\nabla_{Y}X-[X,Y]$.
Convention 2.1.
In the paper we shall use the capital Latin letters $X,Y,\ldots$ to denote smooth vector fields on a smooth manifold $M$, which commute, $[X,Y]=0$. Hence, $T(X,Y)=\nabla_{X}Y-\nabla_{Y}X$.
Let $\nabla^{g}$ denote the Levi-Civita connection corresponding to the symmetric non-degenerate (0,2)-tensor $g$. The J.-L. Koszul’s formula reads
$2g(\nabla^{g}_{X}Y,Z)=Xg(Y,Z)+Yg(X,Z)-Zg(Y,X).$
Denote the torsion (0,3)-tensor with respect to $g$ by the same letter, $T(X,Y,Z):=g(T(X,Y),Z)$.
A linear connection $\nabla$ with a torsion tensor $T$ on a generalized Riemannian manifold $(M,G)$ is completely determined by the torsion tensor and the covariant derivative $\nabla g$ of the symmetric part $g$ of $G$, see [12, Theorem 2.2]. Using the vector fields defined in Convention 2.1, the co-boundary formula for exterior derivative of a 2-form $F$ reduces to the following:
$\,dF(X,Y,Z)=X(F(Y,Z))+Y(F(Z,X))+Z(F(X,Y)).$
Here this formula is without the coefficient 3, unlike [5].
Definition 2.2.
Let $(M,G=g+F)$ be a generalized Riemannian manifold and $A\colon TM\to TM$ a skew-symmetric endomorphism of constant rank, see (2.1). A linear connection $\nabla$ on $M$ is said to have the A-torsion condition if its torsion tensor $T$ satisfies the following:
(2.2) $T(AX,Y)=T(X,AY)\quad\text{for all }X,Y\in\mathfrak{X}_{M}.$
2.2. Codazzi connections
A linear connection $\nabla$ on a generalized Riemannian manifold $(M,G=g+F)$ is called a Codazzi connection if the covariant derivative of the symmetric part of the metric tensor $G$ satisfies the following equation, e.g., [2, 22]:
$(\nabla_{Z}\,g)(X,Y)=(\nabla_{X}\,g)(Z,Y),\quad\forall X,Y,Z\in\mathfrak{X}_{M}.$
Such a pair $(G,\nabla)$ is called a Codazzi structure on $(M,G)$.
As a consequence, the tensor $\nabla g$ is completely symmetric in all three variables, that is,
(2.3) $(\nabla_{Z}\,g)(X,Y)=(\nabla_{X}\,g)(Z,Y)=(\nabla_{Y}\,g)(X,Z).$
This structure plays a fundamental role in information geometry and the theory of statistical connections, enabling a deep relationship between the geometric properties of the metric and the statistical properties of models.
Theorem 2.3.
Let $(M,G=g+F)$ be a generalized Riemannian manifold equipped with a Codazzi connection $\nabla$ having a torsion tensor $T$. Then $\nabla$ is uniquely determined by the following formula:
$2g(\nabla_{X}Y,Z)=2g(\nabla^{g}_{X}Y,Z)+T(X,Y,Z)-(\nabla_{X}\,g)(Y,Z).$
The covariant derivative $\nabla F$ (of the skew-symmetric part $F$) of $G$ is given by
$2(\nabla_{X}F)(Y,Z)$ $=2(\nabla^{g}_{X}F)(Y,Z)+T(X,Y,AZ)+T(Z,X,AY)$
$+T(AZ,X,Y)+T(AZ,Y,X)+T(X,AY,Z)+T(Z,AY,X)$
$+(\nabla_{Z}\,g)(AY,X)-(\nabla_{Y}\,g)(AZ,X).$
Proof.
Using (2.3), it holds
$(\nabla_{X}\,g)(AY,Z)=(\nabla_{AY}\,g)(Z,X),\quad(\nabla_{X}\,g)(Y,AZ)=(\nabla_{AZ}\,g)(Y,X).$
Then, by applying [12, Theorem 2.2] the result follows directly. ∎
Example 2.4.
In the sense of Eisenhart’s definition, see [8], a generalized Riemannian manifold is a differentiable manifold equipped with a non-symmetric basic tensor $G(X,Y)=g(X,Y)+F(X,Y)$, with a linear connection explicitly defined by the equality
(2.4) $g(\nabla_{X}Y,Z)=\frac{1}{2}\big{[}XG(Y,Z)+YG(Z,X)-ZG(Y,X)\big{]}=g(\nabla^{g}_{X}Y,Z)+\dfrac{1}{2}dF(X,Y,Z),$
Using the definition of the connection (2.4), we have the following:
(2.5) $g(\nabla_{X}Y,Z)+g(Y,\nabla_{X}Z)$ $=Xg(Y,Z).$
By using (2.5), the covariant derivative of the metric $g$ is the following:
$(\nabla_{X}g)(Y,Z)$ $=X(g(Y,Z))-g(\nabla_{X}Y,Z)-(Y,\nabla_{X}Z)$
$=X(g(Y,Z))-X(g(Y,Z))=0.$
Thus, the metric $g$ is covariantly constant with respect to $\nabla$. The torsion (0,3)-tensor $T$ is totally skew-symmetric and is given by the equation $T(X,Y,Z)=dF(X,Y,Z)$. Therefore, the connection on a generalized Riemannian manifold in the sense of Eisenhart’s definition is a Codazzi connection.
2.3. Generalized metric connections
Here, we consider generalized metric connections, i.e., linear connections $\nabla$ preserving the generalized Riemannian metric, $G$, on a generalized Riemannian manifold $(M,G=g+F)$. First, we recall the following.
Theorem 2.5 (see [12]).
Let $(M,G=g+F)$ be a generalized Riemannian manifold.
  1. i)
    A linear connection $\nabla$ on $M$ preserves the generalized Riemannian metric $G$ if and only if it preserves its symmetric part $g$ and its skew-symmetric part $F$: $\nabla G=0\Leftrightarrow\nabla g=\nabla F=0\Leftrightarrow\nabla g=\nabla A=0.$
  2. ii)
    If $\nabla$ is a generalized metric connection with a torsion tensor $T$, then the following holds:
    $2(\nabla^{g}_{X}F)(Y,Z)=$ $-T(X,Y,AZ)-T(Z,X,AY)$
    (2.6) $-T(AZ,X,Y)-T(AZ,Y,X)-T(X,AY,Z)-T(Z,AY,X);$
    in particular, the exterior derivative of $F$ satisfies the following equality:
    (2.7) $dF(X,Y,Z)=-T(X,Y,AZ)-T(Y,Z,AX)-T(Z,X,AY);$
    Conversely, if the condition (ii)) is valid, then there exists a unique generalized metric connection $\nabla$ on $(M,G=g+F)$ determined by the torsion $T$ by the following formula:
    (2.8) $2g(\nabla_{X}Y,Z)=2g(\nabla^{g}_{X}Y,Z)+T(X,Y,Z)+T(Z,X,Y)-T(Y,Z,X).$
Remark 2.6.
According to i) of Theorem 2.5, a generalized Riemannian manifold $(M,G=g+F)$ equipped with a generalized metric connection $\nabla$ can be seen as a Riemann-Cartan manifold $(M,g,\nabla)$ equipped with the fundamental 2-form $F$.
The formulas in Theorem 2.5 are especially meaningful under the assumption that the torsion tensor $T$ satisfies the $A$-torsion condition (2.2) with totally skew-symmetric property.
Proposition 2.7.
Let $(M,G=g+F)$ be a generalized Riemannian manifold with non-degenerate fundamental 2-form $F$. Then a generalized metric connection $\nabla$ with a totally skew-symmetric torsion (0,3)-tensor $T$ satisfies the $A$-torsion condition (2.2) if and only if $(\nabla^{g}_{AX}A)X=0\ (X\in\mathfrak{X}_{M})$.
Proof.
Using (ii)) and totally skew-symmetric torsion (0,3)-tensor, by replacing $X$ with $AX$ in the first term and $Y$ with $AY$ in the second term, we obtain
$(\nabla^{g}_{AX}F)(Y,Z)+(\nabla^{g}_{AY}F)(X,Z)=-\dfrac{1}{2}\big{[}T(AX,Y,AZ)-T(X,AY,AZ)\big{]}.$
Since $A$ is a non-singular endomorphism, from the above the assertion follows. ∎
Definition 2.8.
Let $\nabla$ be a linear connection on a generalized Riemannian manifold and $\nabla^{g}$ the Levi-Civita connection of $g$. The contorsion (or, difference) (1,2)-tensor $K$ is defined by
$K(X,Y)=\nabla_{X}Y-\nabla^{g}_{X}Y.$
The contorsion (0,3)-tensor is defined by $K(X,Y,Z):=g(K(X,Y),Z)$.
The Levi-Civita connection is a distinguished point in affine space of all linear connections $\nabla$ on $M$ that is in one-to-one correspondence with the linear space of all contorsion tensors $K$ on $(M,g)$.
By Theorem 2.5, any generalized metric connection on a generalized Riemannian manifold is completely determined by the torsion tensor $T$. If $\nabla$ is a generalized metric connection on $(M,G=g+F)$, then from (2.8) we have
$2K(X,Y,Z)=T(X,Y,Z)+T(Z,X,Y)-T(Y,Z,X).$
In this case, the contorsion (0,3)-tensor $K(X,Y,Z)$ is totally skew-symmetric if and only if the torsion (0,3)-tensor $T(X,Y,Z)$ is totally skew-symmetric.
2.4. The Nijenhuis tensor
The Nijenhuis tensor $N_{P}$ of a (1,1)-tensor $P$ on a smooth manifold $M$ is defined as, e.g.[20],
(2.9) $N_{P}(X,Y)=[PX,PY]+P^{2}[X,Y]-P[PX,Y]-P[X,PY].$
The Nijenhuis tensor is skew-symmetric by definition and it plays a fundamental role in almost complex (resp. almost para-complex) geometry. We denote the Nijenhuis (0,3)-tensor with respect to $g$ with the same letter, $N_{P}(X,Y,Z):=g(N_{P}(X,Y),Z)$.
If $A^{2}=-{\rm Id}$ (resp. $A^{2}={\rm Id}$) then the celebrated Nulander-Nirenberg theorem (see, e.g. [20]) shows that an almost complex structure is integrable if and only if the Nijenhuis tensor $N_{A}$ vanishes.
Let $\nabla$ be a linear connection on $M$ preserving the generalized Riemannian metric $G,\ \nabla G=0$. Then $\nabla$ preserves $g,F$ and $A$$\nabla g=\nabla F=\nabla A=0$. Using the definition of the torsion tensor $T$ and the covariant derivative $\nabla A$, we express the Nijenhuis tensor in terms of $T$ and $\nabla A$ as follows:
$N_{A}(X,Y)$ $=(\nabla_{AX}A)Y-(\nabla_{AY}A)X-A(\nabla_{X}A)Y+A(\nabla_{Y}A)X$
(2.10) $-T(AX,AY)-A^{2}T(X,Y)+AT(AX,Y)+AT(X,AY).$
Using $\nabla A=0$ in (2.4) and applying (2.1), we get the following:
(2.11) $N_{A}(X,Y,Z)=-T(AX,AY,Z)-T(X,Y,A^{2}Z)-T(AX,Y,AZ)-T(X,AY,AZ).$
2.5. The totally skew-symmetric torsion (0,3)-tensor
Pseudo-Riemannian manifolds equipped with metric connections having a totally skew-symmetric torsion (0,3)-tensor $T$, i.e. the following condition is true:
(2.12) $T(X,Y,Z)=-T(X,Z,Y),$
became very attractive in the last twenty years mainly due to the relations with supersymmetric string theories (see e.g. [27, 16, 3] and references therein, for a mathematical treatment consult the nice overview [1]). The main point is that the number of preserved supersymmetries is equal to the number of parallel spinors with respect to such a connection.
Remark 2.9.
If the torsion (0,3)-tensor is totally skew-symmetric, (2.12), then the $Q$-torsion condition (3.1) follows from the $A$-torsion condition (2.2), but the converse does not hold in general.
For a generalized Riemannian manifold $(M,G=g+F)$ equipped with a generalized metric connection $\nabla$ that has a totally skew-symmetric torsion (0,3)-tensor, the following is true:
(2.13) $N_{A}(X,Y,AZ)+N_{A}(X,Z,AY)=dF(X,Y,A^{2}Z)+dF(X,Z,A^{2}Y),$
see Theorem  2.5 in [12]. Moreover, the torsion tensor $T$ of $\nabla$ satisfies the following equalities:
(2.14) $\begin{split}T(AX,AY,Z)=-N_{A}(X,Y,Z)+dF(X,Y,AZ),\\ T(AX,Y,Z)=2(\nabla^{g}_{X}F)(Y,Z)-dF(X,Y,Z),\end{split}$
and is determined by the following formula:
(2.15) $T(X,Y,Z)=2g(\nabla_{X}Y,Z)-2g(\nabla^{g}_{X}Y,Z).$
If, in addition, the fundamental 2-form $F$ of the generalized Riemannian metric $G$ is closed, $dF=0$, then (2.13) and (2.14) hold inserting $dF=0$.
3. New Applications of Weak Metric Structures
A generalized Riemannian metric $G$ is equivalent to a choice of a (pseudo-)Riemannian metric $g$ and a 2-form $F$ (a (1,1)-tensor $A$ satisfying (2.1)), such that $G=g+F$. A generalized metric connection, i.e. a linear connection preserving $G$, is a metric connection preserving the fundamental 2-form $F$, or equivalently, a Riemannian connection preserving the (1,1)-tensor $A$. This supplies a number of examples (given below using weak metric structures defined by V. Rovenski and R. Wolak in [24]) with a (1,1)-tensor $A$ of constant rank. Weak almost contact structures generalize the classical almost contact structures, i.e., the linear complex structure on the contact distribution is replaced by a nonsingular skew-symmetric (1,1)-tensor, hence the identity operator in the structural equations is replaced by a nonsingular self-adjoint (1,1)-tensor $Q$.
For $\nabla Q=0$ and the Nijenhuis (0,3)-tensor $N_{Q}(X,Y,Z):=g(N_{Q}(X,Y),Z)$, we have
$N_{Q}(X,Y,Z)=-T(QX,QY,Z)-T(X,Y,Q^{2}Z)+T(QX,Y,QZ)+T(X,QY,QZ).$
Definition 3.1.
Let $(M,G=g+F)$ be a generalized Riemannian manifold. A linear connection $\nabla$ is said to have $Q$-torsion condition if its torsion tensor $T$ satisfies
(3.1) $T(QX,Y)=T(X,QY)=QT(X,Y)\quad\text{for all }X,Y\in\mathfrak{X}_{M},$
where $Q:TM\to TM$ is an endomorphism that is self-adjoint with respect to the symmetric part $g$ of metric $G$, i.e., $g(QX,Y)=g(X,QY)$ for all $X,Y\in\mathfrak{X}_{M}$.
Note that the $Q$-torsion condition is trivial when $Q=\lambda\,{\rm Id}$ for some positive real $\lambda$ (or in the classical case). The $A$-torsion condition is not trivial when $Q=\lambda\,{\rm Id}$.
3.1. Weak almost Hermitian structure
Let us consider a weak almost Hermitian manifold $(M,g,A,Q)$, i.e. a Riemannian manifold $(M,g)$ of dimension $n\,(=2m\geq 4)$ endowed with non-singular endomorphisms: $A$ (skew-symmetric) and $Q$ (self-adjoint) and the fundamental 2-form $F$, such that the following conditions hold, see [23]:
$A^{2}=-Q,\quad g(AX,AY)=g(QX,Y),\quad F(X,Y)=g(AX,Y).$
From $A^{2}=-Q$ we conclude that $A$ commutes with $Q$: $[A,Q]=0$; hence $F(X,QY)=F(QX,Y).$
Definition 3.2.
A weak almost Hermitian manifold is said to be weak nearly Kähler if the covariant derivative of $A$ with respect to the Levi-Civita connection $\nabla^{g}$ is skew-symmetric:
$(\nabla^{g}_{X}A)X=0\quad\Longleftrightarrow\quad(\nabla^{g}_{X}F)(X,Y)=0.$
If $\nabla^{g}A=0$, then such $(M,g,A,Q)$ is called a weak Kähler manifold.
At the same time, $(M,g,A,Q)$ admits a generalized Riemannian structure $G=g+F$. Suppose that a linear connection $\nabla$ on $(M,G=g+F)$ preserves $G$, then it also preserves the metric $g$, the form $F$, and the tensor $A$, see Theorem 2.5, and we have $\nabla g=\nabla F=\nabla A=0$. Consequently, the connection $\nabla$ also preserves the tensor $Q$, i.e., $\nabla Q=0$. Thus, using the contorsion tensor $K$, (2.15) and its relation with the torsion tensor, $K(X,Y,Z):=g(K(X,Y),Z)=\frac{1}{2}T(X,Y,Z)$, we get
(3.2) $(\nabla^{g}_{X}A)Y=\nabla^{g}_{X}(AY){-}A\nabla^{g}_{X}Y=AK(X,Y){-}K(X,AY)=\dfrac{1}{2}\big{[}AT(X,Y){-}T(X,AY)\big{]},$
(3.3) $(\nabla^{g}_{X}Q)Y=QK(X,Y)-K(X,QY)=\dfrac{1}{2}\big{[}QT(X,Y)-T(X,QY)\big{]}.$
For the case of totally skew-symmetric torsion (0,3)-tensor of $\nabla$, the $A$-torsion condition (2.2) gives
(3.4) $T(X,AY)=T(AX,Y)=-AT(X,Y),\quad T(QX,Y)=T(X,QY)=QT(X,Y),$
(3.5) $T(X,QY)=-T(AX,AY)=AT(X,AY)=AT(AX,Y)=T(QX,Y)=QT(X,Y).$
From (3.2) we conclude that the $A$-torsion condition implies
(3.6) $(\nabla^{g}_{X}A)Y=-AT(X,Y).$
This leads us to the following theorem.
Theorem 3.3.
Let $(M,g,A,Q)$ be a weak almost Hermitian manifold with a fundamental 2-form $F$, considered as a generalized Riemannian manifold $(M,G=g+F)$. Suppose that $\nabla$ is a generalized metric connection ((((i.e., $\nabla G=0)$ with totally skew-symmetric torsion (0,3)-tensor $T$. Then $\nabla$ satisfies the $A$-torsion condition (2.2) if and only if $(M,g,A,Q)$ is a weak nearly Kähler manifold. In the case of (2.2), the torsion tensor of $\nabla$ is determined by the structure tensors as follows:
(3.7) $\begin{split}T(AX,Y,Z)&=-\dfrac{1}{3}\,dF(X,Y,Z),\\ T(QX,Y,Z)&=\dfrac{1}{4}\,N_{A}(X,Y,Z)\ \big{(}=\dfrac{1}{3}\,dF(AX,Y,Z)\ \big{)}.\end{split}$
Proof.
Assume that the connection $\nabla$ with totally skew-symmetric torsion (0,3)-tensor satisfies the $A$-torsion condition (2.2). From (3.6) with $Y=X$ we conclude that $(M,g,A,Q)$ is a weak nearly Kähler manifold. Conversely, assuming $(\nabla^{g}_{X}F)(X,Z)=0$ and taking into account (ii)), we obtain $T(AX,Y,Z)=T(X,AY,Z)$, which implies that the linear connection $\nabla$ satisfies the $A$-torsion condition. Further, using (3.4) in (2.7), we obtain
$T(AX,Y,Z)=-\dfrac{1}{3}\,dF(X,Y,Z).$
Next, using (3.5) in (2.11) and (2.14), we find that $N_{A}$ is totally skew-symmetric and
$T(QX,Y,Z)=\dfrac{1}{4}N_{A}(X,Y,Z),$
that completes the proof of (3.7). ∎
Theorem 3.4.
Let the conditions of Theorem 3.3 be satisfied. Then the $Q$-torsion condition (3.1)
(i) is equivalent to the following: the Levi-Civita connection preserves the tensor $Q$, i.e., $\nabla^{g}\,Q=0$.
(ii) provides the following: the Nijenhuis tensor $N_{Q}$, see (2.9) with $O=Q$, vanishes, thus $Q$ is integrable.
Proof.
(i) Using (3.3), we have $2\,(\nabla^{g}_{X}Q)(Y)=QT(X,Y)-T(X,QY)$. Therefore, $\nabla^{g}Q=0$ if and only if $QT(X,Y)=T(X,QY)$, which is exactly the $Q$-torsion condition in the case of totally skew-symmetric (0,3)-tensor $T$ and self-adjoint $Q$:
$g(QT(X,Y),Z)$ $=g(T(X,Y),QZ)=T(X,Y,QZ)=T(Y,QZ,X)=g(T(Y,QZ),X)$
$\quad=g(T(QY,Z),X)=T(QY,Z,X)=T(X,QY,Z)=g(T(X,QY),Z).$
This completes the proof of (i).
(ii) Let the totally skew-symmetric torsion (0,3)-tensor satisfies the $Q$-torsion condition. So, we have
$T(QX,Y,Z)=T(X,QY,Z)=T(X,Y,QZ).$
Then
$T(QX,QY,Z)$ $=T(X,Q^{2}Y,Z)=T(X,Y,Q^{2}Z),$
$T(QX,Y,QZ)$ $=T(X,QY,QZ)=T(X,Y,Q^{2}Z).$
Substituting the last two equalities into $N_{Q}$, we have
$N_{Q}(X,Y,Z)=-T(X,Y,Q^{2}Z)-T(X,Y,Q^{2}Z)+T(X,Y,Q^{2}Z)+T(X,Y,Q^{2}Z)=0.$
This completes the proof of (ii). ∎
Example 3.5.
The Chern connection $\nabla^{C}$, see [7], referred to in some papers as the canonical or Hermitian connection is the unique linear connection on a Hermitian manifold $(M,g,A)$ preserving the generalized Riemannian metric $G=g+F$, so that the torsion tensor $T$ of $\nabla^{C}$ has the property $T(X,AY)=T(AX,Y)$ for all $X,Y\in\mathfrak{X}_{M}$. This, implies
$T(AX,Y,Z)=T(X,AY,Z)=-T(X,Y,AZ)\quad\text{for all }X,Y,Z\in\mathfrak{X}_{M}.$
The connection $\nabla^{C}$ does not have totally skew-symmetric torsion (0,3)-tensor, and its torsion tensor satisfies different properties than the $A$-torsion condition (with an opposite sign in the term $AT(X,Y)$ compared to (3.4)). This provides motivation for introducing the $A$-torsion condition (2.2) and extending $\nabla^{C}$ to the weak Hermitian structure $(g,A)$ on $M$. Using $A^{2}=-Q$, we get
$T(AX,AY,Z)=-T(QX,Y,Z)=-T(X,QY,Z)=-T(X,Y,QZ)\quad\text{for all }X,Y,Z\in\mathfrak{X}_{M}.$
The torsion tensor of $\nabla^{C}$ is determined by
$T(QX,Y,Z)=-\dfrac{1}{2}\big{[}dF(AX,Y,Z)+dF(X,AY,Z)\big{]}$
and the following is true: $T(AX,Y,Z)=(\nabla^{g}_{Z}F)(X,Y)$.
Example 3.6.
The Bismut connection $\nabla^{B}$, see [4], also known as Strominger or Strominger-Bismut connection, is the unique linear connection with totally skew-symmetric torsion (0,3)-tensor on a Hermitian manifold $(M,g,A)$ preserving the generalized Riemannian metric $G=g+F$. This provides motivation for extending $\nabla^{B}$ to the weak Hermitian structure $(g,A)$. Using $A^{2}=-Q$ and $N_{A}=0$, we have
$\begin{split}T(QX,Y,Z)&=T(X,QY,Z)=T(X,Y,QZ)\\ &=T(AX,AY,Z)+T(AX,Y,AZ)+T(X,AY,AZ)\quad\text{for all }X,Y,Z\in\mathfrak{X}_{M}.\end{split}$
From (2.14), the torsion (0,3)-tensor of $\nabla^{B}$ is determined by
$T(AX,AY,Z)=dF(X,Y,AZ)$
and the following is true: $dF(QX,Y,Z)=dF(X,QY,Z)=dF(X,Y,QZ)$. In general, on an almost weak Hermitian manifold, the connection $\nabla^{B}$ has no totally skew-symmetric torsion (0,3)-tensor.
The results obtained here provide a natural generalization of the previously established results for almost Hermitian (and almost para-Hermitian) manifolds [10, 14]. Our approach thus extends and unifies these results within more general “weak” geometric structures.
The 6-dimensional unit sphere $\mathbb{S}^{6}$ in the set of all purely imaginary Cayley numbers is a classical example of a non-Kähler nearly Kähler manifold. Other examples of 6-dimensional non-Kähler nearly Kähler manifolds include $\mathbb{S}^{3}\times\mathbb{S}^{3}$, the complex projective space $\mathbb{CP}^{3}$ and the flag manifold $\mathbb{F}^{3}$.
The following example presents many non-Kähler weak Kähler manifolds as well as non-nearly Kähler weak nearly Kähler manifolds.
Example 3.7.
L. P. Eisenhart [9] proved that if a Riemannian manifold $(M,g)$ admits a parallel symmetric 2-tensor other than the constant multiple of $g$, then it is reducible. Therefore, a weak Kähler manifold $(M,g,A,Q)$ with $Q\neq\lambda\,{\rm Id}$ for some $\lambda\in C^{\infty}(M)$ is reducible.
Take two (or more) almost Hermitian manifolds $(M_{j},g_{j},J_{j})$, thus $J_{j}^{2}=-{\rm Id}_{\,j}$. The product $\prod_{\,j}(M_{j},g_{j},\sqrt{\lambda_{j}}\,J_{j})$, where $\lambda_{j}\neq 1$ are different positive constants, is a weak almost Hermitian manifold with $Q=\bigoplus_{\,j}\lambda_{j}\,{\rm Id}_{\,j}$. We call $\prod_{\,j}(M_{j},g_{j},\sqrt{\lambda_{j}}\,J_{j})$ a $(\lambda_{1},\ldots,\lambda_{k})$-weighed product of almost Hermitian manifolds, in this case $(M,g=\oplus_{\,j}\,g_{j})$ decomposes into $k$ weak almost Hermitian factors.
Moreover, if $(M_{j},g_{j},A_{j})$ are (nearly) Kähler manifolds, then their $(\lambda_{1},\ldots,\lambda_{k})$-weighed product is a non-(nearly) Kähler weak (nearly) Kähler manifold. According to the above, any non-Kähler weak Kähler manifold is locally a weighed product of Kähler manifolds. The classification of weak nearly Kähler manifolds in dimensions $\geq 4$ is an open problem. Some 4-dimensional non-Kähler weak Kähler manifolds appear as $(\lambda_{1},\lambda_{2})$-weighed products of 2-dimensional Kähler manifolds. Some 6-dimensional non-Kähler weak Kähler manifolds appear as $(\lambda_{1},\lambda_{2},\lambda_{3})$-weighed products of 2-dimensional Kähler manifolds or $(\lambda_{1},\lambda_{2})$-weighed products of 2-dimensional and 4-dimensional Kähler manifolds, and similarly for even dimensions $>6$. Some 8-dimensional non-nearly Kähler weak nearly Kähler manifolds appear as $(\lambda_{1},\lambda_{2})$-weighed products of 2-dimensional Kähler manifolds and 6-dimensional nearly Kähler manifolds, and similarly for even dimensions $>8$.
Such $(\lambda_{1},\ldots,\lambda_{k})$-weighed products of almost Hermitian manifolds can serve as new models of generalized Riemannian manifolds for theoretical physics.
Theorem 3.8.
Let conditions of Theorem 3.3 be satisfied.
(i) Suppose that $Q=\lambda\,{\rm Id}$ for some $\lambda\in C^{\infty}(M)$. Then $\lambda=const>0$, the $Q$-torsion condition (3.1) is true, and $(\lambda^{-1/2}A,g)$ is an almost Hermitian structure; moreover, if the linear connection $\nabla$ satisfies the $A$-torsion condition (2.2), then $(\lambda^{-1/2}A,g)$ is a nearly Kähler structure.
(ii) Suppose that $Q\neq\lambda\,{\rm Id}$ for any $\lambda\in C^{\infty}(M)$. Then there exist $k>1$ mutually orthogonal even-dimensional distributions ${\mathcal{D}}_{i}\subset TM\ (1\leq i\leq k)$ such that $\bigoplus_{\,i}{\mathcal{D}}_{i}=TM$ and ${\mathcal{D}}_{i}$ are the eigen-distributions of $Q$ with different constant eigenvalues $\lambda_{i}:$ $0<\lambda_{1}<\ldots<\lambda_{k}$; moreover, if the $Q$-torsion condition (3.1) ((((or the $A$-torsion condition (2.2))))) is true, then the distributions ${\mathcal{D}}_{i}$ are involutive and define $\nabla^{g}$-totally geodesic foliations ${\mathcal{F}}_{i}$ and $M^{2m+s}(A,Q,g)$ is locally the $(\lambda_{1},\ldots,\lambda_{k})$-weighed product of almost Hermitian manifolds ((((nearly Kähler manifolds, respectively)))).
Proof.
(i) Since $\nabla Q=0$, we get $\lambda=const>0$, hence $(\lambda^{-1/2}A,g)$ is an almost Hermitian structure. If the $A$-torsion condition is true, then by Theorem 3.3, $(\lambda^{-1/2}A,g)$ is a nearly Kähler structure.
(ii) Suppose that $Q=-A^{2}$ is not conformal. For a weak almost Hermitian structure $(g,A)$ on $M^{2m}$ we will build an $A$-basis at a point $x\in M$ consisting of mutually orthogonal nonzero vectors of $T_{x}M$. Let $e_{1}\in T_{x}M$ be a unit eigenvector of the self-adjoint operator $Q>0$ with the minimal eigenvalue $\lambda_{1}\neq 0$. Then, $Ae_{1}\in T_{x}M$ is orthogonal to $e_{1}$ and $Q(Ae_{1})=A(Qe_{1})=\lambda_{1}Ae_{1}$. Thus, the subspace of $T_{x}M$ orthogonal to the plane Span$\{e_{1},Ae_{1}\}$ is $Q$-invariant (and $A$-invariant). Note that $g(e_{1},e_{1})=g(Ae_{1},Ae_{1})=g(Qe_{1},e_{1})=\lambda_{1}$. Continuing in the same manner, we find a basis $\{e_{1},Ae_{1},\ldots,e_{m},Ae_{m}\}$ of $T_{x}M$ consisting of mutually orthogonal vectors. Hence, $Q$ has $k$ different eigenvalues $0<\lambda_{1}<\ldots<\lambda_{k}$ of even multiplicities $n_{1},\ldots,n_{k}$, and $\sum\nolimits_{\,i=1}^{k}n_{i}=2m$. In this basis, $A$ and $Q$ have block-diagonal structures: $Q=[\lambda_{1}{\rm Id}_{\,n_{1}},\ldots,\lambda_{k}{\rm Id}_{\,n_{k}}]$ and $A=[\sqrt{\lambda_{1}}J_{\,n_{1}},\ldots,\sqrt{\lambda_{k}}J_{\,n_{k}}]$, where $J_{\,n_{i}}$ is a complex structure on a $n_{i}$-dimensional subspace of $T_{x}M$.
Since $Q$ is $\nabla$-parallel, we get the same structure at every point of $M$, that is, $k$ and all $\lambda_{i}$ are constant on $M$, and there exist mutually orthogonal $\nabla$-parallel (and $A$-invariant) eigen-distributions ${\mathcal{D}}_{i}$ of $Q$ with constant different eigenvalues $\lambda_{i}$. Let the $Q$-torsion condition (3.1) be true, then using (3.1) for any vector fields $X,Y\in{\mathcal{D}}_{i}$, i.e., $QX=\lambda_{i}X$ and $QY=\lambda_{i}Y$, we have
$Q[X,Y]$ $=Q\{\nabla_{X}Y-\nabla_{Y}X-T(X,Y)\}=\nabla_{X}(QY)-\nabla_{Y}(QX)-QT(X,Y)$
$=\lambda_{i}\{\nabla_{X}Y-\nabla_{Y}X\}-\lambda_{i}T(X,Y)=\lambda_{i}[X,Y].$
Hence each distribution ${\mathcal{D}}_{i}$ is involutive and defines a foliation ${\mathcal{F}}_{i}$. Similarly we can show that $Q(\nabla_{X}Y)=\nabla_{X}(QY)=\lambda_{i}\nabla_{X}Y$, that is, each ${\mathcal{F}}_{i}$ is a $\nabla$-totally geodesic foliation. Then (by Theorem 3.4) $\nabla^{g}Q=0$, hence each ${\mathcal{F}}_{i}$ is a $\nabla^{g}$-totally geodesic foliation and by de Rham Decomposition Theorem (see [20]), $(M,g)$ splits into $(\lambda_{1},\ldots,\lambda_{k})$-weighed product of almost Hermitian manifolds. Now, let the $A$-torsion condition (2.2) be true, then by Theorem 3.3, $(M,g)$ splits into the $(\lambda_{1},\ldots,\lambda_{k})$-weighed product of nearly Kähler manifolds. ∎
3.2. Weak almost contact metric structure
Contact Riemannian geometry is of growing interest due to its important role in both theoretical physics and pure mathematics. Weak almost contact structures, i.e., the complex structure on the contact distribution is approximated by a non-singular skew-symmetric tensor, allowed us to take a new look at the theory of contact manifolds and find new applications. A weak almost contact metric manifold $M^{2m+1}(A,Q,\eta,\xi,g)$ is a $(2m+1)$-dimensional Riemannian manifold equipped with a (1,1)-tensor $A$ of rank $2m$, a non-singular (1,1)-tensor field $Q\neq{\rm Id}$, a 1-form $\eta$ and a vector field $\xi$ dual to $\eta$ with respect to the metric $g$$\eta(\xi)=1,\ \eta(X)=g(X,\xi)$, satisfying the following compatibility conditions:
(3.8) $A^{2}=-Q+\eta\otimes\xi,\quad g(AX,AY)=g(QX,Y)-\eta(X)\eta(Y),\quad A\xi=0,\quad Q\xi=\xi.$
In this case, the skew-symmetric part $F$, $F(X,Y)=g(AX,Y)$, of $G=g+F$ is degenerate, $F(\xi,X)=0$, and has rank $2m$. Denote by ${\mathcal{D}}=A(TM)$ a contact distribution on $M$.
Since $\nabla A=0$, the distribution ${\mathcal{D}}$ is $\nabla$-parallel; hence its unit normal, $\xi$, is also $\nabla$-parallel: $\nabla\xi=0$. By duality, $\eta=g(\xi,\cdot)$, we conclude that the 1-form $\eta$ is also $\nabla$-parallel: $\nabla\eta=0$.
On the other side, taking $\nabla$-derivative of $A\xi=0$ and using $\nabla A=0$, we obtain $A\nabla_{X}\,\xi=0$. Hence $\nabla_{X}\,\xi\,||\,\xi$ for any $X$. On the other hand, taking derivative of $g(\xi,\xi)=1$ and using $\nabla g=0$, we obtain $g(\nabla_{X}\xi,\xi)=0$. Therefore, $\nabla_{X}\xi=0$ for any $X$, i.e., $\nabla\xi=0$.
Then (3.8) (see also Theorem 2.5) gives $\nabla Q=0$:
00 0 $=(\nabla_{X}A)AY+A(\nabla_{X}A)Y=(\nabla_{X}A^{2})Y$
$=-(\nabla_{X}Q)Y+(\nabla_{X}\,\eta)(Y)\,\xi-\eta(Y)\nabla_{X}\,\xi=-(\nabla_{X}Q)Y.$
Since $T$ is totally skew-symmetric, the equation (2.15) yields $0=g(\nabla^{g}_{X}\,\xi,Y)+\dfrac{1}{2}T(X,\xi,Y),$ which shows that $\xi$ is a geodesic and Killing vector field: $\nabla^{g}_{\xi}\,\xi=0$ and $g(\nabla^{g}_{X}\,\xi,Y)+g(\nabla^{g}_{Y}\,\xi,X)=0$. Further, we have
(3.9) $d\eta(X,Y)=(\nabla^{g}_{X}\,\eta)(Y)-(\nabla^{g}_{Y}\,\eta)X)=g(\nabla^{g}_{X}\,\xi,Y)-g(\nabla^{g}_{Y}\,\xi,X)=T(X,Y,\xi)$
and
$d\eta(X,\xi)=-d\eta(\xi,X)=0.$
Now, using (3.8) in equation (2.11), the Nijenhuis tensor becomes:
(3.10) $N_{A}(X,Y,Z)+\eta(Z)d\eta(X,Y)=T(X,Y,QZ)-T(AX,AY,Z)-T(AX,Y,AZ)-T(X,AY,AZ).$
The right-hand side of (3.10) is totally skew-symmetric if and only if the torsion tensor satisfies the $Q$-torsion condition (3.1): $T(X,Y,QZ)=T(X,QY,Z)$. Let us introduce the (0,3)-tensor
(3.11) $N_{A}^{\mathrm{wac}}=N_{A}+d\eta\otimes\eta,$
which we will call the Nijenhuis tensor in the weak almost contact geometry, similarly to the definition given by Blair (see, e.g., [5], where this formula is with the coefficient 2) in the case of almost contact structures. From (3.10) and (3.9) and taking into account (3.8), we obtain
$N_{A}(X,Y,\xi)=-d\eta(AX,AY),\quad N_{A}(\xi,Y,Z)=d\eta(Y,QZ)-d\eta(AY,AZ).$
So, the last equations give
$\begin{split}d\eta(Y,QZ)=N_{A}(\xi,Y,Z)-N_{A}(Y,Z,\xi).\end{split}$
The above considerations yield the following theorem.
Theorem 3.9.
Let $M^{2m+1}(A,Q,\eta,\xi,g)$ be a weak almost contact metric manifold considered as a generalized Riemannian manifold $(M,G=g+F)$. Suppose that $\nabla$ is a generalized metric connection ((((i.e., $\nabla G=0)$ with totally skew-symmetric torsion (0,3)-tensor $T$. Then the following properties hold:
  1. i)
    The Reeb vector field $\xi$ is a geodesic and Killing vector field.
  2. ii)
    The connection $\nabla$ satisfies the $Q$-torsion condition (3.1) if and only if the Nijenhuis tensor in weak almost contact geometry $N^{\mathrm{wac}}$ given by  (3.11) is totally skew-symmetric.
The following theorem is similar to Theorem 3.8.
Theorem 3.10.
Let conditions of Theorem 3.9 be satisfied. Then the following properties hold.
(i) If $Q|_{\mathcal{D}}=\lambda\,{\rm Id}_{\mathcal{D}}$ for some $\lambda\in C^{\infty}(M)$, then $\lambda=const$ and $(\lambda^{-1/2}A,\eta,\xi,g)$ is an almost contact metric structure on $M^{2m+1}$; moreover, if the connection $\nabla$ satisfies the $Q$-torsion condition (3.1) ((((or the $A$-torsion condition (2.2))))), then $M^{2m+1}(\lambda^{-1/2}A,\eta,\xi,g)$ is locally the product of $\mathbb{R}$ and an almost Hermitian manifold ((((a nearly Kähler manifold, respectively)))).
(ii) If $Q|_{\mathcal{D}}\neq\lambda\,{\rm Id}_{\mathcal{D}}$ for any $\lambda\in C^{\infty}(M)$, then there exist $k>1$ mutually orthogonal even-dimensional distributions ${\mathcal{D}}_{i}\subset{\mathcal{D}}$ such that $\bigoplus_{\,i=1}^{k}{\mathcal{D}}_{i}={\mathcal{D}}$ and ${\mathcal{D}}_{i}$ are eigen-distributions of $Q$ with constant eigenvalues $\lambda_{i}:$ $0<\lambda_{1}<\ldots<\lambda_{k}$; moreover, if the $Q$-torsion condition (3.1) ((((or the $A$-torsion condition (2.2))))) is true, then the distributions ${\mathcal{D}}_{i}$ are involutive and define $\nabla^{g}$-totally geodesic foliations ${\mathcal{F}}_{i}$ and $M^{2m+1}(A,Q,\eta,\xi,g)$ is locally a $(1,\lambda_{1},\ldots,\lambda_{k})$-weighed product of $\mathbb{R}$ and $k$ almost Hermitian manifolds ((((nearly Kähler manifolds, respectively)))).
Proof.
(i) Since $\nabla Q=0$, we get $\lambda=const>0$, hence $(\lambda^{-1/2}A,\eta,\xi,g)$ is an almost contact metric structure. If the connection $\nabla$ satisfies the $A$-torsion condition (2.2), then by Theorem 3.3, $M^{2m+1}(\lambda^{-1/2}A,\eta,\xi,g)$ is locally the product of $\mathbb{R}$ and a nearly Kähler manifold.
(ii) Suppose that $Q|_{\mathcal{D}}$ is not conformal. For a weak almost contact metric manifold, as well as for a weak almost Hermitian manifold, see the proof of Theorem 3.8, there exists an $A$-basis $\{e_{1},Ae_{1},\ldots,e_{m},Ae_{m};\xi\}$ of $T_{x}M$ consisting of mutually orthogonal nonzero vectors. Hence, $Q$ restricted on ${\mathcal{D}}_{x}$ has $k>1$ different eigenvalues $0<\lambda_{1}<\ldots<\lambda_{k}$ of even multiplicities $n_{1},\ldots,n_{k}$, and $\sum\nolimits_{\,i=1}^{k}n_{i}=2m$. In this basis, $A$ and $Q$ restricted on ${\mathcal{D}}_{x}$ have block-diagonal structures: $Q|_{\mathcal{D}}=[\lambda_{1}{\rm Id}_{n_{1}},\ldots,\lambda_{k}{\rm Id}_{n_{k}}]$ and $A|_{\mathcal{D}}=[\sqrt{\lambda_{1}}J_{n_{1}},\ldots,\sqrt{\lambda_{k}}J_{n_{k}}]$, where $J_{n_{i}}$ is a complex structure on a $n_{i}$-dimensional subspace of ${\mathcal{D}}_{x}$. Since $Q$ is $\nabla$-parallel, we get the same structure at each point of $M$, i.e., $k$ and all $\lambda_{i}$ are constant on $M$, and there exist mutually orthogonal $\nabla$-parallel (and $A$-invariant) eigen-distributions ${\mathcal{D}}_{i}\subset{\mathcal{D}}$ of $Q$ with constant different eigenvalues $\lambda_{i}$. The rest of the proof is similar to the proof of Theorem 3.8. ∎
Remark 3.11.
Condition $\nabla G=0$ (hence $\nabla g=\nabla A=0$) in Theorem 3.10 is sufficiently strong. When the contact distribution ${\mathcal{D}}=A(TM)$ has constant dimension $<\dim M$, it makes sense to examine the weaker condition: $\nabla G=0$ along ${\mathcal{D}}$.
Remark 3.12 (Weak $f$-structure).
An $f$-structure by K. Yano [29], where $f^{3}+f=0$, on a smooth manifold $M^{2m+s}$ serves as a higher-dimensional analog of almost complex structures ($s=0$) and almost contact structures ($s=1$). The following generalizes the $f$-structure. Let us consider a weak $f$-manifold $(M,g,A,Q)$, i.e., a (pseudo-)Riemannian manifold $(M,g)$ of dimension $n=2m+s$ endowed with a (1,1)-tensor $f=A$ of rank $2m$, and a non-singular (1,1)-tensor $Q$, satisfying the following compatibility conditions, see [24, 26]:
(3.12) $A^{3}+AQ=0,\quad g(AX,A^{2}Y)=g(QX,AY),\quad F(X,Y)=g(AX,Y).$
Note that if $A$ is non-degenerate ($s=0$), then $A^{3}+AQ=0$ reduces to $A^{2}=-Q$, see Section 3.1.
By the first equality in (3.12), $Q$ commutes with $A$: $[A,Q]=0$. Due to (3.12), we have complementary orthogonal distributions ${\mathcal{D}}=A(TM)$ and ${\mathcal{D}}_{0}=\ker A$, where $\dim{\mathcal{D}}=2m$ and $\dim{\mathcal{D}}_{0}=s$.
We omit here our result with $s>1$, for a weak $f$-manifold with a fundamental 2-form $F$, considered as a generalized Riemannian manifold $(M^{2m+s},G=g+F)$, similar to Theorem 3.10.
3.3. Weak para-Hermitian structure and weak para-$f$-structure
A weak almost para-Hermitian manifold $(M,g,A,Q)$ is a (pseudo-) Riemannian manifold $(M,g)$ of dimension $n\,(=2m\geq 4)$ endowed with non-singular endomorphisms: $A$ (skew-symmetric) and $Q\neq{\rm Id}$ (self-adjoint) and the fundamental 2-form $F$, such that the following conditions hold, see [23]:
$A^{2}=Q,\quad g(AX,AY)=-g(QX,Y),\quad F(X,Y)=g(AX,Y).$
The results analogous to Theorems 3.3 and 3.8 from Section 3.1 also hold for the class of weak para-Hermitian manifolds, and their statements and proofs can be applied here accordingly.
The difference in signs appears in formula (3.7), where, under the conditions of Theorem 3.3, the Nijenhuis tensor is totally skew-symmetric and satisfies
$T(QX,Y,Z)=-\dfrac{1}{4}N_{A}(X,Y,Z)\ \big{(}=-\dfrac{1}{3}dF(AX,Y,Z)\ \big{)}.$
The following generalizes the classical para-$f$-structure, where $f^{3}-f=0$. A weak para-$f$-manifold $(M,g,A,Q)$ is a pseudo-Riemannian manifold $(M,g)$ of dimension $n\,(=2m+s)$ endowed with a (1,1)-tensor $f=A$ of rank $2m$, and a non-singular (1,1)-tensor $Q\neq{\rm Id}$, satisfying the following compatibility conditions, see [25]:
(3.13) $A^{3}-AQ=0,\quad g(AX,A^{2}Y)=-g(QX,AY),\quad F(X,Y)=g(AX,Y).$
Note that if $A$ is non-degenerate ($s=0$), then $A^{3}-AQ=0$ reduces to $A^{2}=Q$. Then ${\mathcal{D}}=A(TM)$ and ${\mathcal{D}}_{0}=\ker A$ are two complementary orthogonal distributions on $M$. We have ${\mathcal{D}}={\mathcal{D}}^{+}\oplus{\mathcal{D}}^{-}$, where ${\mathcal{D}}^{+}$ and ${\mathcal{D}}^{-}$ are positive and negative eigen-distributions of $Q$ restricted to ${\mathcal{D}}$, with $\dim{\mathcal{D}}^{+}=\dim{\mathcal{D}}^{-}=m$ and $\dim{\mathcal{D}}_{0}=s$. By the first equality in (3.13), $Q$ commutes with $A$: $[A,Q]=0$. Weak metric para-$f$-manifolds are introduced and can be examined similarly to weak metric $f$-manifolds.
We pay special attention to weak almost para-contact metric manifolds $M^{2m+1}(A,Q,\eta,\xi,g)$, i.e. a $(2m+1)$-dimensional pseudo-Riemannian manifold of signature $(m+1,m)$ equipped with a (1,1)-tensor $A$ of rank $2m$, a 1-form $\eta$, a vector field $\xi$ dual to $\eta$ with respect to the metric $g$$\eta(\xi)=1,\ \eta(X)=g(X,\xi)$, and a non-singular (1,1)-tensor $Q$, satisfying the following conditions:
(3.14) $A^{2}=Q-\eta\otimes\xi,\ \ g(AX,AY)=-g(QX,Y)+\eta(X)\eta(Y),\ \ A\xi=0,\ \ Q\xi=\xi.$
In this case, the skew-symmetric part $F$ of $G=g+F$ has rank equal to $2m$ and $F(\xi,X)=0$ holds.
In a similar way, for the totally skew-symmetric torsion (0,3)-tensor $T$ of a metric connection $\nabla$ on $M$, we conclude that $\xi$ is both a geodesic and a Killing vector field and $d\eta(X,Y)=T(X,Y,\xi)$ holds. By applying (3.14) in (2.11), the Nijenhuis tensor of $A$ becomes:
(3.15) $N_{A}(X,Y,Z){-}\eta(Z)d\eta(X,Y)=-T(X,Y,QZ){-}T(AX,AY,Z){-}T(AX,Y,AZ){-}T(X,AY,AZ).$
The right-hand side of (3.15) is totally skew-symmetric if and only if the torsion tensor of $\nabla$ satisfies the $Q$-torsion condition (3.1). In the same manner, we introduce the (0,3)-tensor
(3.16) $N_{A}^{\mathrm{wapc}}=N_{A}-d\eta\otimes\eta,$
which we will call the Nijenhuis tensor in weak almost para-contact geometry.
From (3.15) and $d\eta(X,Y)=T(X,Y,\xi)$, taking into account (3.14), we obtain
$N_{A}(X,Y,\xi)=-d\eta(AX,AY),\quad N_{A}(\xi,Y,Z)=d\eta(Y,QZ)-d\eta(AY,AZ).$
Thus, the following theorems (similar to Theorems 3.9 and 3.10) hold.
Theorem 3.13.
Let $M^{2m+1}(A,Q,\eta,\xi,g)$ be a weak almost para-contact metric manifold considered as a generalized Riemannian manifold $(M,G=g+F)$. Suppose that there exists a generalized metric connection $\nabla$, i.e., $\nabla G=0$, with a totally skew-symmetric torsion (0,3)-tensor. Then the following hold:
  1. i)
    the Reeb vector field $\xi$ is a geodesic and Killing vector field.
  2. ii)
    the connection $\nabla$ satisfies the $Q$-torsion condition (3.1) if and only if the Nijenhuis tensor in weak almost para-contact geometry $N^{\mathrm{wapc}}$ given by  (3.16) is totally skew-symmetric.
Theorem 3.14.
Let conditions of Theorem 3.13 be satisfied. Then the following properties hold.
(i) If $Q|_{\mathcal{D}}=\lambda\,{\rm Id}_{\mathcal{D}}$ for some $\lambda\in C^{\infty}(M)$, then $\lambda=const$ and $(\lambda^{-1/2}A,\eta,\xi,g)$ is an almost para contact metric structure on $M^{2m+1}$; moreover, if the connection $\nabla$ satisfies the $Q$-torsion condition (3.1) ((((or the $A$-torsion condition (2.2))))), then $M^{2m+1}(\lambda^{-1/2}A,\eta,\xi,g)$ is locally the product of $\mathbb{R}$ and an almost para-Hermitian manifold ((((a nearly para-Kähler manifold, respectively)))).
(ii) If $Q|_{\mathcal{D}}\neq\lambda\,{\rm Id}_{\mathcal{D}}$ for any $\lambda\in C^{\infty}(M)$, then there exist $k>1$ mutually orthogonal even-dimensional distributions ${\mathcal{D}}_{i}\subset{\mathcal{D}}$ such that $\bigoplus_{\,i=1}^{k}{\mathcal{D}}_{i}={\mathcal{D}}$ and ${\mathcal{D}}_{i}$ are eigen-distributions of $Q$ with constant eigenvalues $\lambda_{i}:$ $0<\lambda_{1}<\ldots<\lambda_{k}$; moreover, if the $Q$-torsion condition (3.1) ((((or the $A$-torsion condition (2.2))))) is true, then the distributions ${\mathcal{D}}_{i}$ are involutive and define $\nabla^{g}$-totally geodesic foliations ${\mathcal{F}}_{i}$ and $M^{2m+1}(A,Q,\eta,\xi,g)$ is locally a $(1,\lambda_{1},\ldots,\lambda_{k})$-weighed product of $\mathbb{R}$ and $k$ almost para-Hermitian manifolds ((((nearly para-Kähler manifolds, respectively)))).
4. Conclusion
This paper presents new applications of weak contact metric structures to generalized Riemannian manifolds equipped with a metric connection of totally skew-symmetric torsion tensor. We described the behavior of the Nijenhuis tensor for a skew-symmetric non-singular endomorphism $A$ and a symmetric non-singular endomorphism $Q$ with respect to symmetric part of metric in specific cases. The vanishing of the Nijenhuis tensor $N_{Q}$, ensured by the $Q$-torsion condition, indicates that $Q$ defines a Nijenhuis operator. This naturally leads to the consideration of Nijenhuis manifolds (e.g. [6]). We have shown that weighted products of almost Hermitian manifolds can serve as new models of generalized Riemannian manifolds that can be useful for theoretical physics. Following the idea, in the next paper we will consider weak structures on the generalized Riemannian manifolds satisfying the Einstein metricity condition [12]. It would also be interesting to characterize generalized Riemannian manifolds equipped with weak metric structures of various distinguished classes, e.g. weak (almost) ${\mathcal{K}}$-, weak (almost) ${\mathcal{C}}$-, weak (almost) ${\mathcal{S}}$-, and weak $f$-${\mathcal{K}}$-contact structures, see [26], in terms of the totally skew-symmetric torsion tensor of a metric connection.
Acknowledgments. M.Z. was partially supported by the Ministry of Education, Science and Technological Development of the Republic of Serbia, project no. 451-03-137/2025-03/200124 and by the Bulgarian Ministry of Education and Science, Scientific Programme ”Enhancing the Research Capacity in Mathematical Sciences (PIKOM)”, No. DO1-67/05.05.2022.